# The Honeycomb Conjecture in normed planes and an alpha-convex variant of a theorem of Dowker

### v2

@@ 4. Results about not necessarily polygonal norms
First, observe that by our conditions, any supporting line of $K$ intersects $K$ in a unique point, and for any point $p$ of $\operatorname{bd}(K)$ there is a unique supporting line of $K$ that contains $p$. This yields, in particular, that for any $n\geq 3$ and any minimum area convex $n$-gon $Q_{n}$ circumscribed about $K$, $Q_{n}$ has exactly $n$ sides, and the tangent points on those sides are distinct. For any point $p\in\operatorname{bd}(K)$, we denote the unique closed half plane supporting $K$ at $p$ by $H(p)$, and the boundary of $H(p)$ by $L(p)$. We call boundary points of $K$ with opposite unit normal vectors opposite, or antipodal points. Note that if $p,q\in\operatorname{bd}(K)$ are not antipodal points of $K$, there there is a unique closed arc of $\operatorname{bd}(K)$ connecting them that does not contain antipodal points. We denote this arc by $\widehat{pq}$, and denote by $A(\widehat{pq})$ the area of the bounded, connected region $R(\widehat{pq})$ whose boundary consists of $\widehat{pq}$ and the segments of $L(p)$ and $L(q)$ connecting $p,q$ to the intersection point of the two lines. We observe that the problem of finding $A_{K}(n)$ coincides with the problem of finding a tiling of $\operatorname{bd}(K)$ by $n$ arcs $\widehat{x_{i}x_{i+1}}$, $i=1,2,n$, $x_{n+1}=x_{1}$ such that $f=\sum_{i=1}^{n}A(\widehat{x_{i},x_{i+1}})$ is minimal.
First, we fix an arbitrary integer $n\geq 4$, and show that $A_{K}(n-1)+A_{K}(n+1)>2A_{K}(n)$. To do it, we follow the proof of the original theorem of Dowker [8, Theorem 1]
Let $Q_{n-1}$ and $Q_{n+1}$ denote a minimum area convex $(n-1)$-gon and $(n+1)$-gon, respectively, circumscribed about $K$. Let the tangent points of $K$ on the sides of $Q_{n-1}$ be $p_{1},p_{2},\ldots,p_{n-1}$ in counterclockwise order, and we denote the tangent points of $Q_{n+1}$ in counterclockwise order on $\operatorname{bd}(K)$ by $q_{1},q_{2},\ldots,q_{n+1}$, where the indices are understood mod $(n-1)$ and $(n+1)$, respectively. Note that then $Q_{n-1}=\bigcap_{i=1}^{n-1}H(p_{i})$ and $Q_{n+1}=\bigcap_{j=1}^{n+1}H(q_{j})$. Since the areas of these regions are finite, for any two consecutive tangent points $p_{i}$, $p_{i+1}$, no arc $\widehat{p_{i}p_{i+1}}$ contains antipodal points of $K$, and the same is true for the points $q_{j}$.
Due to the possible existence of coinciding points in the above two sequences, we unite these sequences as a single sequence $v_{1},v_{2},\ldots,v_{2n}$ in such a way that the points are in this counterclockwise order in $\operatorname{bd}(K)$, $v_{1}=x_{1}$, and removing all the $p_{i}$ (resp. all the $q_{j}$) from this sequence we obtain the sequence $q_{1},\ldots,q_{n+1}$ (resp. $p_{1},\ldots,p_{n-1}$). In the proof we regard this sequence as a cyclic sequence, where the indices are determined mod $2n$, and, with a little abuse of notation, we say that $\widehat{v_{i}v_{j}}$ covers $\widehat{v_{k}v_{l}}$ if $\widehat{v_{k}v_{l}}\subseteq\widehat{v_{i}v_{j}}$ and $i<k<l<j<i+2n$.
@@
Note that since $\mathcal{T}_{0}$ is the union of two tilings consisting of $(n-1)$ and $(n+1)$ arcs, respectively, $\mathcal{T}_{0}$ contains covering pairs. Assume that $\widehat{v_{i}v_{j}}$ covers $\widehat{v_{k}v_{l}}$. Then let $\mathcal{T}_{1}$ denote the $2$-tiling of $\operatorname{bd}(K)$ in which $\widehat{v_{i}v_{j}}$ and $\widehat{v_{k}v_{l}}$ are replaced by $\widehat{v_{i}v_{l}}$ and $\widehat{v_{k}v_{j}}$. Our main observation, as in the proof of [8, Theorem 1] is that in this case
(4.1) $A(\widehat{v_{i}v_{j}})+A(\widehat{v_{k}v_{l}})\geq A(\widehat{v_{i}v_{l}})+A(\widehat{v_{k}v_{j}}),$
@@
According to our conditions, $\sum_{S\in\mathcal{T}_{0}}A(S)\geq\sum_{S\in\mathcal{T}_{1}}A(S)$. Furthermore, as $\mathcal{T}_{0}$ and $\mathcal{T}_{1}$ are $2$-tilings, and $\widehat{v_{k}v_{l}}$ is already covered twice by the arcs considered in the modification, the number of covering pairs in $\mathcal{T}_{1}$ is strictly less than in $\mathcal{T}_{0}$. Repeating this procedure we obtain a $2$-tiling $\mathcal{T}_{m}$ of $\operatorname{bd}(K)$ for which $\sum_{S\in\mathcal{T}_{0}}A(S)\geq\sum_{S\in\mathcal{T}_{,}}A(S)$ and which does not contain covering pairs. Then, $\mathcal{T}_{m}$ decomposes into two tilings $\{\widehat{v_{1},v_{3}},\widehat{v_{3}v_{5}},\ldots,\widehat{v_{2n-1}v_{1}}\}$ and $\{\widehat{v_{2},v_{4}},\widehat{v_{4}v_{6}},\ldots,\widehat{v_{2n}v_{2}}\}$, each of which contains exactly $n$ arcs.
@@
Without loss of generality, let $p_{2}=q_{2}$ corresponding to $v_{3}$ and $v_{4}$, respectively, in the united sequence of vertices. Then $\{p_{1},q_{1}\}=\{v_{1},v_{2}\}$ and $\{p_{3},q_{3}\}=\{v_{5},v_{6}\}$. For $i=1,3$, let $p_{i}^{\prime}$ denote the intersection point of $L(p_{i})$ and $L(p_{2})$, and let $q_{i}^{\prime}$ denote the intersection point of $L(q_{i})$ and $L(q_{2})=L(p_{2})$. Note that since $Q_{n-1}$ is a minimum area convex $(n-1)$-gon circumscribed about $K$, the tangent point on every side is the midpoint of the side. Thus, $p_{2}$ is the midpoint of $[p_{1}^{\prime},p_{3}^{\prime}]$, and we obtain similarly that $q_{2}=p_{2}$ is the midpoint of $[q_{1}^{\prime},q_{3}^{\prime}]$.
Case 1, $p_{1}=v_{1}$, $p_{3}=v_{6}$. Then, since $V_{1},V_{2}$ are minimum area circumscribed $n$-gons, we have that $p_{2}=q_{2}$ is the midpoint of $[p_{1}^{\prime},q_{3}^{\prime}]$ as well as that of $[q_{1}^{\prime},p_{3}^{\prime}]$. But then $||p_{3}^{\prime}-p_{2}||=||q_{3}^{\prime}-p_{2}||=||p_{1}^{\prime}-p_{2}||=||q_{3}^{\prime}-p_{2}||$, implying that $p_{1}=q_{1}$ and $p_{3}=q_{3}$.
@@
Until now, we have shown that for any $n\geq 4$, we have $A_{K}(n-1)+A_{K}(n+1)>2A_{K}(n)$. Thus, for any $n\geq 4$, either $A^{\bar{\alpha}}_{K}(n-1)+A^{\bar{\alpha}}_{K}(n+1)>2A^{\bar{\alpha}}_{K}(n)$ is satisfied for all $\bar{\alpha}>0$, or there is a value $10<\alpha_{n}<$ 1 such that $A^{\bar{\alpha}}_{K}(n-1)+A^{\bar{\alpha}}_{K}(n+1)\geq 2A^{\bar{\alpha}}_{K}(n)$ is satisfied if and only if $\bar{\alpha}\geq\alpha_{n}$.
@@ 5. A conjecture of Steinhaus
[…]
@@
Let $\mathcal{F}$ denote the family of convex disks with unit perimeter and containing $o$. Then for any $K\in\mathcal{F}$, $\operatorname{area}(K)\geq\frac{1}{\lambda}$ and $K\subset\mathbf{B}^{2}$, and thus, $\mathcal{F}$ is universally bounded. Since $\operatorname{cr}(K)\leq 1$ for every $K\in\mathcal{F}$, we need to show that there is some $\tau>0$ such that $\operatorname{ir}(K)\geq\tau$ for all $K\in\mathcal{F}$. Suppose for contradiction that this is not so. Then, there is a sequence $\{K_{n}\}$ of elements of $\mathcal{F}$ such that $\operatorname{ir}(K_{n})\to 0$. It is known (see e.g. p. 215, Ex. 6.2 of [27]) that for any convex disk we have $\operatorname{ir}(K)\geq\frac{w(K)}{3}$, where $w(K)$ is the minimal width of $K$. Thus, we have $w(K_{n})\to 0$. But this, combined with $K_{n}\subset\mathbf{B}^{2}$, clearly contradicts our assumption that $\operatorname{area}(K_{n})\geq\frac{1}{\lambda}$ for all values of $n$. ∎
@@
Step B: Let $\bigcup\mathcal{F}_{k-1}=N_{k-1}\mathbf{S}$. We fill the region obtained as the closure of $(N_{k-1}+2)\mathbf{S}\setminus(N_{k-1}\mathbf{S})$ by $44N_{k-1}+$ 4 translates of $\mathbf{S}$.
@@
[…]
Fix any positive number $\nu>0$. First, consider the case that we apply Step A $k_{1}$ times, and after that we apply Step B once. Then $N_{k_{1}}=3^{k_{1}}$, and thus, $\mathcal{T}_{k_{1}}$ has $4\cdot 3^{k_{1}}$ boundary edges. In addition, $\mathcal{T}_{k_{1}}$ consists of $14k_{1}+$ 1 cells. We count the number of interior edges: Note that $\mathcal{T}_{0}$ has $4$ edges. Then in each step before the $k_{1}$th one we add $12$ edges, which are all interior edges of $\mathcal{T}_{k_{1}}$. Finally, in the $k_{1}$th step we add four interior edges of $\mathcal{T}_{k_{1}}$. Thus, in total, $\mathcal{T}_{k_{1}}$ has $412k_{1}-$ 4 interior edges. Now, if we set $R_{1}=N_{k_{1}}+\frac{1}{2}$, then
@@ 7. Acknowledgments
The authors thank Frank Morgan for a valuable observation, stated as Remark 2.8 in the paper.

### v4

@@ 4. Results about not necessarily polygonal norms
[…]
First, we fix an arbitrary integer $n\geq 4$, and show that $A_{K}(n-1)+A_{K}(n+1)>2A_{K}(n)$. To do it, we follow the proof of the original theorem of Dowker [8, Theorem 1].
Let $Q_{n-1}$ and $Q_{n+1}$ denote a minimum area convex (n1)(n-1)-gon and $(n+1)$-gon, respectively, circumscribed about $K$. Let the tangent points of $K$ on the sides of $Q_{n-1}$ be $p_{1},p_{2},\ldots,p_{n-1}$ in counterclockwise order, and we denote the tangent points of $Q_{n+1}$ in counterclockwise order on $\operatorname{bd}(K)$ by $q_{1},q_{2},\ldots,q_{n+1}$, where the indices are understood mod (n1)(n-1) and $(n+1)$, respectively. Note that then $Q_{n-1}=\bigcap_{i=1}^{n-1}H(p_{i})$ and $Q_{n+1}=\bigcap_{j=1}^{n+1}H(q_{j})$. Since the areas of these regions are finite, for any two consecutive tangent points $p_{i}$, $p_{i+1}$, no arc $\widehat{p_{i}p_{i+1}}$ contains antipodal points of $K$, and the same is true for the points $q_{j}$.
Due to the possible existence of coinciding points in the above two sequences, we unite these sequences as a single sequence $v_{1},v_{2},\ldots,v_{2n}$ in such a way that the points are in this counterclockwise order in $\operatorname{bd}(K)$, $v_{1}=p_{1}$. In the proof we regard this sequence as a cyclic sequence, where the indices are determined mod $2n$, and, with a little abuse of notation, we say that $\widehat{v_{i}v_{j}}$ covers $\widehat{v_{k}v_{l}}$ if $\widehat{v_{k}v_{l}}\subseteq\widehat{v_{i}v_{j}}$ and $i<k<l<j<i+2n$.
@@
Note that since $\mathcal{T}_{0}$ is the union of two tilings consisting of (n1)(n-1) and $(n+1)$ arcs, respectively, $\mathcal{T}_{0}$ contains covering pairs. Assume that $\widehat{v_{i}v_{j}}$ covers $\widehat{v_{k}v_{l}}$. Then let $\mathcal{T}_{1}$ denote the $2$-tiling of $\operatorname{bd}(K)$ in which $\widehat{v_{i}v_{j}}$ and $\widehat{v_{k}v_{l}}$ are replaced by $\widehat{v_{i}v_{l}}$ and $\widehat{v_{k}v_{j}}$. Our main observation, as in the proof of [8, Theorem 1] is that in this case
[…]
@@
According to our conditions, $\sum_{S\in\mathcal{T}_{0}}A(S)\geq\sum_{S\in\mathcal{T}_{1}}A(S)$. Furthermore, as $\mathcal{T}_{0}$ and $\mathcal{T}_{1}$ are $2$-tilings, and $\widehat{v_{k}v_{l}}$ is already covered twice by the arcs considered in the modification, the number of covering pairs in $\mathcal{T}_{1}$ is strictly less than in $\mathcal{T}_{0}$. Repeating this procedure we obtain a $2$-tiling $\mathcal{T}_{m}$ of $\operatorname{bd}(K)$ for which $\sum_{S\in\mathcal{T}_{0}}A(S)\geq\sum_{S\in\mathcal{T}_{,}}A(S)$ and which does not contain covering pairs. Then, $\mathcal{T}_{m}$ decomposes into two tilings $\{\widehat{v_{1}v_{3}},\widehat{v_{3}v_{5}},\ldots,\widehat{v_{2n-1}v_{1}}\}$ and $\{\widehat{v_{2}v_{4}},\widehat{v_{4}v_{6}},\ldots,\widehat{v_{2n}v_{2}}\}$, each of which contains exactly $n$ arcs.
@@
Without loss of generality, let $p_{2}=q_{2}$ corresponding to $v_{3}$ and $v_{4}$, respectively, in the united sequence of vertices. Then $\{p_{1},q_{1}\}=\{v_{1},v_{2}\}$ and $\{p_{3},q_{3}\}=\{v_{5},v_{6}\}$. For $i=1,3$, let $p_{i}^{\prime}$ denote the intersection point of $L(p_{i})$ and $L(p_{2})$, and let $q_{i}^{\prime}$ denote the intersection point of $L(q_{i})$ and $L(q_{2})=L(p_{2})$. Note that since $Q_{n-1}$ is a minimum area convex (n1)(n-1)-gon circumscribed about $K$, the tangent point on every side is the midpoint of the side. Thus, $p_{2}$ is the midpoint of $p_{1}^{\prime},p_{3}^{\prime}]$, and we obtain similarly that $q_{2}=p_{2}$ is the midpoint of $q_{1}^{\prime},q_{3}^{\prime}]$.
Figure 4. The configuration in Case 1 of the proof of Theorem 4.4.
[…]
@@
Until now, we have shown that for any $n\geq 4$, we have $A_{K}(n-1)+A_{K}(n+1)>2A_{K}(n)$. Thus, for any $n\geq 4$, either $A^{\bar{\alpha}}_{K}(n-1)+A^{\bar{\alpha}}_{K}(n+1)>2A^{\bar{\alpha}}_{K}(n)$ is satisfied for all $\bar{\alpha}>0$, or there is a value $0<\alpha_{n}<1$ such that $A^{\bar{\alpha}}_{K}(n-1)+A^{\bar{\alpha}}_{K}(n+1)\geq 2A^{\bar{\alpha}}_{K}(n)$ is satisfied if and only if $\bar{\alpha}\geq\alpha_{n}$.
@@ 5. A conjecture of Steinhaus
If we replace the $\liminf$ in the above definition by $\limsup$, we obtain the upper average isoperimetric ratio $\overline{I}(\mathcal{T})$ of a cell. If these quantities are equal, the common value is called the average isoperimetric ratio of a cell, denoted by $I(\mathcal{T})$.
@@
Let $\mathcal{F}$ denote the family of convex disks, containing $o$, with unit perimeter and isoperimetric ratio at most $\lambda$. Then for any $K\in\mathcal{F}$, $\operatorname{area}(K)\geq\frac{1}{\lambda}$ and $K\subset\mathbf{B}^{2}$, and thus, $\mathcal{F}$ is universally bounded. Since $\operatorname{cr}(K)\leq 1$ for every $K\in\mathcal{F}$, we need to show that there is some $\tau>0$ such that $\operatorname{ir}(K)\geq\tau$ for all $K\in\mathcal{F}$. Suppose for contradiction that this is not so. Then, there is a sequence $\{K_{n}\}$ of elements of $\mathcal{F}$ such that $\operatorname{ir}(K_{n})\to 0$. It is known (see e.g. p. 215, Ex. 6.2 of [27]) that for any convex disk we have $\operatorname{ir}(K)\geq\frac{w(K)}{3}$, where $w(K)$ is the minimal width of $K$. Thus, we have $w(K_{n})\to 0$. But this, combined with $K_{n}\subset\mathbf{B}^{2}$, clearly contradicts our assumption that $\operatorname{area}(K_{n})\geq\frac{1}{\lambda}$ for all values of $n$. ∎
@@
Step B: Let $\bigcup\mathcal{F}_{k-1}=N_{k-1}\mathbf{S}$. We fill the region obtained as the closure of $(N_{k-1}+2)\mathbf{S}\setminus(N_{k-1}\mathbf{S})$ by $4N_{k-1}+4$ translates of $\mathbf{S}$.
@@
In the following part we describe a suitable infinite sequence of $A$s and $B$s, where the kthth element is $A$ or $B$, if $\mathcal{T}_{k}$ is obtained from $\mathcal{T}_{k-1}$ by applying Step A or Step B, respectively. Then we set $\mathcal{T}=\bigcup_{k=1}^{\infty}\mathcal{T}_{k}$. Before describing this sequence, we remark that independently of its definition, if an edge $E_{1}$ of a cell $C_{1}\in\mathcal{T}$ intersects an edge $E_{2}$ of another cell $C_{2}$ in a nondegenerate segment, then $E_{1}\subseteq E_{2}$ or $E_{2}\subseteq E_{1}$. Thus, we can apply an arbitrarily small deformation to the cells of $\mathcal{T}$ to obtain an edge-to-edge convex tiling $\mathcal{T}^{\prime}$ of $\mathbb{R}^{2}$ with the same combinatorial properties; this deformation can be done subsequently after each step. As a consequence, for simplicity, if a side of a polygon, appearing in one of the two steps above, is dissected into smaller segments by the vertices of the neighbors of the cell, we regard each smaller segment as a separate edge of the cell. Using this interpretation, in the following, we call the edges of $\mathcal{T}$ in $\operatorname{bd}\left(\bigcup\mathcal{T}_{k}\right)$ boundary edges of $\mathcal{T}_{k}$, and the edges of $\mathcal{T}$ in $\operatorname{int}\left(\bigcup\mathcal{T}_{k}\right)$ interior edges of $\mathcal{T}_{k}$.
Fix any positive number $\nu>0$. First, consider the case that we apply Step A $k_{1}$ times, and after that we apply Step B once. Then $N_{k_{1}}=3^{k_{1}}$, and thus, $\mathcal{T}_{k_{1}}$ has $4\cdot 3^{k_{1}}$ boundary edges. In addition, $\mathcal{T}_{k_{1}}$ consists of $4k_{1}+1$ cells. We count the number of interior edges: Note that $\mathcal{T}_{0}$ has $4$ edges. Then in each step before the $k_{1}$th one we add $12$ edges, which are all interior edges of $\mathcal{T}_{k_{1}}$. Finally, in the $k_{1}$th step we add four interior edges of $\mathcal{T}_{k_{1}}$. Thus, in total, $\mathcal{T}_{k_{1}}$ has $12k_{1}-4$ interior edges. Now, if we set $R_{1}=N_{k_{1}}+\frac{1}{2}$, then
@@ 7. Acknowledgments
The authors thank Frank Morgan for a valuable observation, stated as Remark 2.8 in the paper, and an unknown referee for many useful observations.